OpenStudy (anonymous):

Which answer shows this equation written in standard form? 9-4(x-y)=-6x+6 A) 2x+4y=-3 B) 2x-4y=3 C) -10x+4y=-3 D) 2x+4y=3

OpenStudy (anonymous):

@Peach-es1234

Vocaloid (vocaloid):

start by distributing -4(x-y)

OpenStudy (anonymous):

Example: Put x2 = 7 into Standard Form Answer: x2 - 7 = 0

OpenStudy (anonymous):

ok...hmm -x-(-y)?

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

not quite hint: A(B-C) = A*B - A*C -4(x-y) = ?

OpenStudy (anonymous):

oh -4*x or -4x -(-4*y) or -4y?

OpenStudy (anonymous):

2x + 4y = –3

OpenStudy (anonymous):

is the answer a?

OpenStudy (anonymous):

oh thx

OpenStudy (anonymous):

So first we solve and simplify pemdas first parenthasees 9-4(x-y)=-6x+6 distribute using distributeive property of multiplication which is a(b+c)=(ab)+(ac) so -4(x-y)=-4x+4y 9-4x+4y=-6x+6 add 6x to both sides 9+2x+4y=6 subtract 9 from both sides 2x+4y=-3 only A matches this A is the correct answer

OpenStudy (anonymous):

ohhh i get it i always hav eproblems with pemdas tho but i get it now!! thx medal for u!

OpenStudy (anonymous):

Your doing great! :) Now another 1 ? :)

OpenStudy (anonymous):

next Q?

OpenStudy (anonymous):

oh ok let me open anoder